You have assigned the following values to these three firms: Price Upcoming Dividend Growth Beta US Bancorp $ 47.45 $ 2.95 6.60 % 1.80 Praxair 60.70 1.80 18.50 2.73 Eastman Kodak 46.25 2.00 7.80 0.75 Assume that the market portfolio will earn 10.80 percent and the risk-free rate is 3.00 percent. Compute the required return for each company using both CAPM and the constant-growth model. (Do not round intermediate calculations. Round your final answer to 2 decimal places.)

CAPM Constant-growth model

US Bancorp required return % %

Praxair required return % %

Eastman Kodak required return % %
Answer #1


Company

CAPM

Constant Growth

Stocks

US Bancorp

17.04%

12.82%

Praxair

24.29%

21.47%

Eastman Kodak

8.85%

12.12%

WORKING:

CAPM (see last column for returns):

Company

Risk free rate

Market Return

Beta

CAPM Working

Expected return

Stocks

Rf

Rm

B

Ri = Rf+B*(Rm-Rf)

Ri = Rf+B*(Rm-Rf)

US Bancorp

3.00%

10.80%

1.80

=3%+1.80*(10.8%-3%)

17.04%

Praxair

3.00%

10.80%

2.73

=3%+2.73*(10.8%-3%)

24.29%

Eastman Kodak

3.00%

10.80%

0.75

=3%+0.75*(10.8%-3%)

8.85%

Constant growth model or DDM (see last row for returns):

Given details

US Bancorp

Praxair

Eastman Kodak

Existing growth rate = g =

6.60%

18.50%

7.80%

Expected dividend = D1 = D0*(1+g) =

2.95

1.80

2.00

Expected rate = r = Cost of equity =

?

?

?

Current stock price = P0 =

47.45

60.70

46.25

Flotation cost = f =

0.00

0.00

0.00

Formula for calculating the Expected rate:

r = (D1/(P0-f))+g =

12.82%

21.47%

12.12%